# Alpha-Z divergence unveils further distinct phenotypic traits of human brain connectivity fingerprint

### v1

@@
Md Kaosar Uddin Department of Mathematics and Statistics,Auburn University, Auburn, Alabama, USA Nghi Nguyen Faculty of Science, Vrije Universiteit Amsterdam, the Netherlands Huajun Huang Department of Mathematics and Statistics,Auburn University, Auburn, Alabama, USA Duy Duong-Tran Department of Mathematics, U.S. Naval Academy, Annapolis, Maryland, USA Department of Biostatistics, Epidemiology and Informatics, University of Pennsylvania, Philadelphia, Pennsylvania,USA Jingyi Zheng Department of Mathematics and Statistics,Auburn University, Auburn, Alabama, USA Correspondence: jzz0121@auburn.edu
@@
The accurate identification of individuals from functional connectomes (FCs) is critical for advancing individualized assessments in neuro/psychiatric research. Traditional methods, such as using Pearson’s correlation, have limitations in capturing the complex, non-Euclidean geometry of FC data, leading to suboptimal performance in identifying individual differences. Recent developments have introduced geodesic distance as a more robust metric; however, its performance is highly sensitive to regularization choices of FCs, which vary by spatial scale and are task-condition specific.To address these challenges, we propose a novel distance measure and divergence-based distance metric, the Alpha-Z Bures-Wasserstein divergence, which provides a more flexible and geometry-aware framework for FC comparison. Unlike prior methods, our approach does not require meticulous parameter tuning and maintains strong identification performance across multiple task conditions and spatial resolutions. We evaluate our approach against both traditional (e.g., Euclidean, Pearson) and state-of-the-art manifold-based distances (e.g., affine-invariant, log-Euclidean, and Bures-Wasserstein) and systematically investigate how varying regularization strengths affect the performance of geodesic distance measures on the Human Connectome Project dataset, spanning multiple cognitive tasks and parcellation resolutions from low to high. Our results demonstrate that the proposed method significantly improves the identification rates over traditional and existing geodesic distance measures, particularly when optimized regularization is applied, and notably in high-dimensional settings where matrix rank deficiencies degrade the performance of existing metrics. Furthermore, we validate the generalizability of our approach across different functional connectivity conditions,including resting-state and task-based fMRI, using multiple parcellation schemes.These findings suggest that the new distance measure and divergence provide a more reliable and generalizable framework for functional connectivity analysis, offering enhanced sensitivity in linking FC patterns to cognitive and behavioral outcomes. This work has the potential to pave the way for more precise and individualized diagnostic tools in clinical neuroscience.
@@ RESULTS
As parcellation resolution increases from 100 to 900 regions, the box-and-whisker distributions in Fig. 4(a) show a clear metric-dependent divergence in fingerprinting performance. Alpha-Z Divergence and Alpha Procrustes show the highest central tendencies across all granularity levels, approaching ceiling levels ($>$ 0.95) by 400–500 parcels, while their inter-quartile ranges contract, signalling consistently strong and stable identification. AI and Log-Euclidean distances track closely at coarse resolutions but plateau near 0.60 after 500 parcels, and their wider boxes and longer whiskers indicate greater variability once dimensionality increases. Bures–Wasserstein remains the least effective throughout, never surpassing the lower-mid accuracy band and showing the broadest dispersion of outcomes.
@@
Fig. 4(b) presents a heat-map of median subject-identification accuracy (ID-rate) for five symmetric–positive-definite distance metrics as the cortical parcellation is progressively refined from 100 to 900 regions. Across all metrics, increasing the number of parcels generally boosts fingerprinting performance, but the magnitude of this gain differs markedly. Alpha-Z dominates at every granularity, climbing steeply from 0.590.590.59 at 100100100 parcels to >0.94>0.94> 0.94 by 400400400 parcels and plateauing at 0.96$\approx$ 0.96≈ 0.96 for 600600600 parcels and above. Alpha Procrustes shows the next-best performance, rising from 0.540.540.54 to 0.650.650.65 across the range, while AI and Log-Euclidean display more modest improvements (peaking around 0.570.570.57). BW is the weakest throughout, never exceeding 0.520.520.52 and flattening near 0.490.490.49 for denser parcellations. These trends indicate that (i) finer parcellation enhances individual identification, but with diminishing returns beyond 600$\approx$ 600≈ 600 parcels, and (ii) the choice of metric is critical, with Alpha-Z offering a pronounced advantage over alternative SPD metrics.
@@
Figure 5: Sensitivity of the Affine-Invariant (AI) distance metric to regularization strength across parcellation levels. The solid line shows identification rates (ID rates) under low regularization conditions ($\tau\leq 1$), where performance rapidly declines as the number of parcels increases. The dash line illustrates the effect of higher regularization values ($\tau>1$), where appropriate tuning improves performance stability across tasks and scales. These findings emphasize the critical role of regularization in geodesic-based metrics and motivate the need for more robust alternatives like which does not depend on regularization parameter $\tau$
@@
[…]
@@
We found support for the latter. For each task, we computed Spearman’s rank-order correlation between the network-level identification rate rankings at coarser granularities and those at granularity 900. We then tested whether these correlations exceeded chance expectations using permutation-based null models, with p-values corrected for the false discovery rate from multiple comparisons. This analysis revealed the granularity levels at which the network rankings began to converge with those at the finest scale (figure 6a). Interestingly, for the resting state and the emotion task, these rankings stabilized as early as granularity 100, where the default mode and control networks already emerged as the most individually distinctive, and the somatomotor and limbic networks as the least. Other tasks required higher resolution: for instance, the motor task only began to exhibit stable rankings at granularity 400, with lower resolutions failing to achieve significant alignment with the rankings observed at 900.
@@
In general, all HCP tasks exhibited stable network-level identification rate rankings starting at granularity 400 and above (figure 6a). These stable rankings align well with expectations derived from task designs. For example, the social task, which demands individualized processing of dynamic social cues, resulted in identification rates in the visual network that are among the highest, which suggests high variability in visual encoding strategies across individuals (figure 6b). In contrast, the language and working memory tasks, which present time-locked blocks of stimuli with predictable perceptual demands, elicited lower-ranked visual network identification rates, consistent with the suppression of intersubject variability in early sensory encoding 43 (figure 6b).
@@ DISCUSSION
In contrast, Alpha-Z Divergence and Alpha Procrustes Distance ( with fixed tuning parameter $\alpha=0.6$ for higher granularity) demonstrate both stability and flexibility . They perform robustly across all tasks and parcellations without requiring extensive regularization, as evidenced by their high identification rates even in the absence of aggressive regularization. Alpha-Z Divergence while choosing fixed tuning parameter ( $\alpha=0.99$ and $\ z=1$), in particular, achieves near-perfect identification rates across all parcellation resolutions, including 900 regions, indicating its ability to handle complex, high-dimensional FC data with minimal performance degradation. This robustness suggests that these metrics are well-suited for large-scale neuroimaging analyses, where the data are often high-dimensional and traditional metrics face difficulties in maintaining accuracy.
@@
Moreover, the visualization reveals significant overlap between FCs from different subjects. Notably from Fig.(8c) for AI, subjects such as 111, 333, and 444 show higher mis-labeling rate that are clustered closely with points from other individuals, indicating that AI Distance has difficulty distinguishing between subjects. The considerable overlap and broad distribution of points suggest that AI fails to capture the unique structural similarities within FCs of the same individual across different sessions, particularly at higher parcellation levels.
In contrast, Fig.(8b), which depicts the performance of Alpha-Z Divergence, demonstrates markedly improved clustering of FCs for the same subject. In the Fig. (8c) for Alpha-Z, the same-subject correct-match percentages are significantly higher across all participants. For instance, subject 111 achieves a similarity rate of 51.0%, a dramatic improvement over the results achieved with AI distance. Similarly, subjects 222 and 555, which exhibited lower labeling rates using AI, show much better performance with Alpha-Z Divergence, achieve correct-match rates of 30.4% and 46.6%, respectively.
@@ ACKNOWLEDGMENTS
Disclaimer
”The views expressed in this manuscript are those of the author(s) and do not reflect the official policy or position of the U.S. Naval Academy, Department of the Navy, the Department of Defense, or the U.S. Government.”
@@
Method details
In this study, we utilized a set of functional brain atlases, specifically the Schaefer parcellation of the cortex. This parcellation is derived from resting-state fMRI data collected from 1,489 participants, which were aligned using surface-based registration techniques. To generate the Schaefer parcellation, a gradient-weighted Markov random field approach was employed, combining local gradient information with global similarity metrics. The Schaefer parcellation is available in ten levels of granularity, ranging from 100 to 1000 regions in increments of 100. These parcellations are provided in both volumetric and grayordinate formats. Since the grayordinate parcellations share the same surface space as the HCP fMRI data, they can be mapped onto the fMRI data with relative ease. Surface-based mapping offers superior alignment between the fMRI data and the Schaefer parcellations compared to volumetric mapping. Therefore, we used surface-based mapping to align the 100–900 region Schaefer parcellations with the fMRI data. During the data processing phase of this study, we were unable to successfully map the 1,000 region Schaefer parcellation for the HCP Young Adult dataset. Additionally, 14 subcortical regions were integrated into each parcellation, as provided by the HCP release filename: Atlas_ROI2.nii.gz.This file was converted from NIFTI to CIFTI format using the HCP Workbench softwareHCP Workbench software, via the command wb_command -cifti-create-label. For example, the Schaefer-100 parcellation resulted in a total of 114 brain regions, and the Schaefer-900 parcellation resulted in a total of 914 brain regions.
In this work, we used data from the HCP 1,200 participants release35 and extracted three different subsets. The first consists of 428 unrelated participants (223 women, mean age: 28.67 years old, range: 22–36) selected.
@@
For task-fMRI data, we followed the same steps but used a broader frequency range for the bandpass filter (0.001–0.250 Hz), given the unclear relationship between different tasks and optimal frequency ranges.
@@ 3. Bures-Wasserstein (BW) Distance
[…]
@@ 4. Alpha-Procrustes Distance
  • =$\mathcal{B}=\{B_{j}\}_{j=1}^{428}$: A set of functional connectivity (FC) matrices representing another session or condition.

### v2

@@
Md Kaosar Uddin Department of Mathematics and Statistics,Auburn University, Auburn, Alabama, USA Nghi Nguyen Faculty of Science, Vrije Universiteit Amsterdam, the Netherlands Huajun Huang Department of Mathematics and Statistics,Auburn University, Auburn, Alabama, USA Duy Duong-Tran Department of Biostatistics, Epidemiology and Informatics, University of Pennsylvania, Philadelphia, Pennsylvania,USA Jingyi Zheng Department of Mathematics and Statistics,Auburn University, Auburn, Alabama, USA Correspondence: jzz0121@auburn.edu
@@
The accurate identification of individuals from functional connectomes (FCs) is critical for advancing individualized assessments in neuro/psychiatric research. Traditional metrics such as Pearson correlation and Euclidean distance fail to capture the non-Euclidean geometry of FCs, while geodesic distances (e.g., affine-invariant, log-Euclidean) require task- and scale-specific regularization and degrade under high-dimensional conditions. To address these challenges, we propose a novel distance measure, the Alpha-Z Bures-Wasserstein divergence, a geometry-aware metric for functional connectome comparison. Unlike prior methods, our approach does not require meticulous parameter tuning and maintains strong identification performance across multiple task conditions, scan lengths, and spatial resolutions. We compare our method with classical (e.g., Euclidean, Pearson) and state-of-the-art manifold-based distances (e.g., affine-invariant, log-Euclidean, Bures–Wasserstein) and assess how changing regularization strengths affects geodesic distance performance across Human Connectome Project tasks and parcellation scales. Our results demonstrate that the proposed method significantly improves the identification rates over traditional and existing geodesic distance measures, particularly when optimized regularization is applied, and notably in high-dimensional settings where matrix rank deficiencies degrade the performance of existing metrics. Furthermore, we validate the generalizability of our approach across different functional connectivity conditions, including resting-state and task-based fMRI, using multiple parcellation schemes. These findings establish Alpha-Z as a reliable and generalizable framework for functional connectivity analysis, enhancing sensitivity to cognitive and behavioral patterns and offering strong potential for individualized clinical neuroscience.
@@ RESULTS
As parcellation resolution increases from 100 to 900 regions, the box-and-whisker distributions in Fig. 4(a) show a clear metric-dependent divergence in fingerprinting performance. Alpha-Z Divergence and Alpha Procrustes show the highest central tendencies across all granularity levels, approaching ceiling levels (>> 0.95) by 400–500 parcels, while their inter-quartile ranges contract, signalling consistently strong and stable identification. AI and Log-Euclidean distances track closely at coarse resolutions but plateau near 0.60 after 500 parcels, and their wider boxes and longer whiskers indicate greater variability once dimensionality increases. Bures–Wasserstein remains the least effective throughout, never surpassing the lower-mid accuracy band and showing the broadest dispersion of outcomes.
@@
Fig. 4(b) presents a heat-map of median subject-identification accuracy (ID-rate) for five symmetric–positive-definite distance metrics as the cortical parcellation is progressively refined from 100 to 900 regions. Across all metrics, increasing the number of parcels generally boosts fingerprinting performance, but the magnitude of this gain differs markedly. Alpha-Z dominates at every granularity, climbing steeply from $0.59$ at $100$ parcels to >0.94>0.94 by $400$ parcels and plateauing at $\approx 0.96$ for $600$ parcels and above. Alpha Procrustes shows the next-best performance, rising from $0.54$ to $0.65$ across the range, while AI and Log-Euclidean display more modest improvements (peaking around $0.57$). BW is the weakest throughout, never exceeding $0.52$ and flattening near $0.49$ for denser parcellations. These trends indicate that (i) finer parcellation enhances individual identification, but with diminishing returns beyond $\approx 600$ parcels, and (ii) the choice of metric is critical, with Alpha-Z offering a pronounced advantage over alternative SPD metrics.
@@
Figure 5: Sensitivity of the Affine-Invariant (AI) distance metric to regularization strength across parcellation levels. The solid line shows identification rates (ID rates) under low regularization conditions ($\tau\leq 1$), where performance rapidly declines as the number of parcels increases. The dash line illustrates the effect of higher regularization values ($\tau>1$), where appropriate tuning improves performance stability across tasks and scales. These findings emphasize the critical role of regularization in geodesic-based metrics and motivate the need for more robust alternatives like which does not depend on regularization parameter $\tau$ .
@@
The performance curves for regularization values ranging from $\tau=10$ to $\tau=48$ show that these higher values stabilize identification rates across most tasks, preventing the sharp declines seen in the low-$\tau$ settings. This stabilization is particularly beneficial for tasks like Social and WM, where the identification rates remain consistent across all parcellations. Furthermore, in tasks like Motor, high regularization values such as $\tau=30$ and $\tau=40$ result in relatively flat performance curves, maintaining identification rates around 0.6 even at higher parcellation resolutions.
@@
We found support for the latter. For each task, we computed Spearman’s rank-order correlation between the network-level identification rate rankings at coarser granularities and those at granularity 900. At this resolution, for all tasks except Emotion, the number of regions within each Yeo network (see table 1) remained lower than the number of time points in the fMRI runs, ensuring that the corresponding FC matrices were not rank-deficient. We then tested whether these correlations exceeded chance expectations using permutation-based null models, with p-values corrected for the false discovery rate from multiple comparisons. This analysis revealed the granularity levels at which the network rankings began to converge with those at the finest scale (figure 6a). Interestingly, for the resting state and the emotion task, these rankings stabilized as early as granularity 100, where the default mode and control networks already emerged as the most individually distinctive, and the somatomotor and limbic networks as the least. Other tasks required higher resolution: for instance, the motor task only began to exhibit stable rankings at granularity 400, with lower resolutions failing to achieve significant alignment with the rankings observed at 900.
@@
In general, all HCP tasks exhibited stable network-level identification rate rankings starting at granularity 400 and above (figure 6a). These stable rankings align well with expectations derived from task designs. For example, the Social task, which demands individualized processing of dynamic social cues, resulted in identification rates in the visual network that are among the highest, which suggests high variability in visual encoding strategies across individuals (figure 6b). In contrast, the Language and Motor tasks, which present time-locked blocks of stimuli with predictable perceptual demands, elicited lower-ranked visual network identification rates, consistent with the suppression of intersubject variability in early sensory encoding 43 (figure 6b).
@@ DISCUSSION
In contrast, Alpha-Z Divergence and Alpha Procrustes Distance ( with fixed tuning parameter $\alpha=0.6$ for higher granularity) demonstrate both stability and flexibility . They perform robustly across all tasks and parcellations without requiring extensive regularization, as evidenced by their high identification rates even in the absence of aggressive regularization. Alpha-Z Divergence while choosing fixed tuning parameter ( $\alpha=0.99$ and z=1\ z=1), in particular, achieves near-perfect identification rates across all parcellation resolutions, including 900 regions, indicating its ability to handle complex, high-dimensional FC data with minimal performance degradation. This robustness suggests that these metrics are well-suited for large-scale neuroimaging analyses, where the data are often high-dimensional and traditional metrics face difficulties in maintaining accuracy.
@@
Moreover, the visualization reveals significant overlap between FCs from different subjects. Notably from Fig.(8c) for AI, subjects such as $1$, 33, and $4$ show higher mis-labeling rate that are clustered closely with points from other individuals, indicating that AI Distance has difficulty distinguishing between subjects. The considerable overlap and broad distribution of points suggest that AI fails to capture the unique structural similarities within FCs of the same individual across different sessions, particularly at higher parcellation levels.
In contrast, Fig.(8b), which depicts the performance of Alpha-Z Divergence, demonstrates markedly improved clustering of FCs for the same subject. In the Fig. (8c) for Alpha-Z, the same-subject correct-match percentages are significantly higher across all participants. For instance, subject $1$ achieves a similarity rate of 51.0%, a dramatic improvement over the results achieved with AI distance. Similarly, subjects $2$ and $5$, which exhibited lower labeling rates using AI, show much better performance with Alpha-Z Divergence, achieve correct-match rates of 30.4% and 46.6%, respectively.
@@ ACKNOWLEDGMENTS
Experimental data
In this study, we utilized a set of functional brain atlases, specifically the Schaefer parcellation of the cortex. This parcellation is derived from resting-state fMRI data collected from 1,489 participants, which were aligned using surface-based registration techniques. To generate the Schaefer parcellation, a gradient-weighted Markov random field approach was employed, combining local gradient information with global similarity metrics. The Schaefer parcellation is available in ten levels of granularity, ranging from 100 to 1000 regions in increments of 100. These parcellations are provided in both volumetric and grayordinate formats. Since the grayordinate parcellations share the same surface space as the HCP fMRI data, they can be mapped onto the fMRI data with relative ease. Surface-based mapping offers superior alignment between the fMRI data and the Schaefer parcellations compared to volumetric mapping. Therefore, we used surface-based mapping to align the 100–900 region Schaefer parcellations with the fMRI data. During the data processing phase of this study, we were unable to successfully map the 1,000 region Schaefer parcellation for the HCP Young Adult dataset. Additionally, 14 subcortical regions were integrated into each parcellation, as provided by the HCP release filename: Atlas_ROI2.nii.gz.This file was converted from NIFTI to CIFTI format using the HCP Workbench softwareHCP Workbench software, via the command wb_command -cifti-create-label. For example, the Schaefer-100 parcellation resulted in a total of 114 brain regions, and the Schaefer-900 parcellation resulted in a total of 914 brain regions. The Schaefer parcellation atlases contain labels of Yeo canonical functional networks 38, 45 whose numbers of regions for all parcellation levels are included in Table 1.
Table 1: Number of ROIs per Yeo network across Schaefer parcellation granularities.
Visual Somatomotor
Dorsal
Attention
Salience /
Ventral
Attention
Limbic Control
Default
Mode
Granularity
100 17 14 15 12 5 13 24
200 29 35 26 22 12 30 46
300 47 57 34 34 20 40 68
400 61 77 46 47 26 52 91
500 74 96 56 59 33 69 113
600 89 112 72 73 42 82 130
700 120 128 87 75 49 92 149
800 134 151 99 87 54 105 170
900 147 173 104 105 60 117 194
In this work, we used data from the HCP 1,200 participants release 46 and extracted three different subsets. The first consists of 428 unrelated participants (223 women, mean age: 28.67 years old, range: 22–36) selected.
@@
FC matrices were constructed by computing the Pearson correlation coefficient between the mean time series of every pair of brain regions. This resulted in symmetric, weighted adjacency matrices with values ranging from –1 to 1. FC matrices were computed for each participant individually.
@@ 3. Bures-Wasserstein (BW) Distance
where $\text{tr}(\cdot)$ denotes the trace of a matrix. The BW distance captures both the spread (variance) and the mean (location) of the distributions, making it a powerful tool for comparing functional connectomes and other brain imaging data where Gaussian assumptions are reasonable.
@@ 4. Alpha-Procrustes Distance
  • $\mathcal{B}=\{B_{j}\}_{j=1}^{428}$: A set of functional connectivity (FC) matrices representing another session or condition.